OpenStudy (anonymous):

The volume of a hexagonal prism is 12 square 3 in3. What is the height of the prism if the base area is 3 square 3 in3? 36 square 3 in 4 square 3 in 4 in 36 in

jhonyy9 (jhonyy9):

do you know what is the formula for volum of a hexagonal prism ?

jhonyy9 (jhonyy9):

so i think this is area of base multiplie height of prism so if the volum is 12sqr3 in^3 and area of base =3sqrt3 in^3 than we can writing that V=A*h from this h=V/A so h=(12sqrt3)/(3sqrt3) =4 h=4 in

OpenStudy (anonymous):

thank you